### v1

@@ 4 Irreducible Subspace Projections
4.1 Representation Theory Preliminaries
Following standard texts on group and representation theory Fulton1991RepresentationTA ; ragone2023representation we provide a brief summary and introduction to the underlying mathematical framework utilised in this section.
@@ 4.1 Representation Theory Preliminaries
If a representation does contains a non-trivial subrepresentation, then it is called a reducible representation. Representations can often be decomposed into a direct sum of their irreducible representations, in which case they are called completely reducible. Maschke’s Theorem maschke_ueber_1898 states that any finite-dimensional representation of a finite group $G$ over a field $F$ will be completely reducible (so long as the field characteristic does not divide the order of the group). It is also the case that every finite-dimensional representation of compact Lie groups is completely reducible hall_lie_2015 . If a representation $(U,V)$ is completely reducible, then there exists a basis in which we can write $U_{g}$ as
@@
4.2 Subspace Projection Circuit
We start by observing a result from the representation theory of finite groups which states serre_linear_reps
@@ 4.2 Subspace Projection Circuit
This is a known result in representation theory; for more details, see reference texts such as Fulton1991RepresentationTA ; serre_linear_reps . In this section, we explore a method for implementing this projection on a quantum circuit using the Linear Combination of Unitaries (LCU) technique. Specifically, we demonstrate the practicality of realising the projection described in Theorem 3 on a quantum device. In the most comprehensive scenario, we establish that a quantum circuit can execute combinations of such $\hat{P}_{r}$ projections at the same time via the LCU method.
@@
Let $\lvert\psi\rangle\in V=(\mathbb{C}^{2})^{\otimes n}$ be a quantum state encoded on $n$ qubits. Let $U:G\rightarrow SU(2^{n})$ be a unitary representation of a compact finite group $G$, where the representatives for group elements $g_{i}\in G$ are denoted $U_{g_{i}}\in SU(2^{n})$ and are unitary operators that acts on $\lvert\psi\rangle$ and can be implemented in a quantum device. Utilising the Linear Combination of Unitaries framework on a quantum circuit, one can probabilistically apply a linear combination of irreducible representation projections
@@
where the $u_{i,j}$ will never affect the algorithm, but are required as the matrices involved must be of size $2^{k}$. The first $R$ columns correspond to the vectors that contain the characters for each element in a given representation. Considering the character orthogonality theorem Huppert1998 which states that
$\frac{1}{\rvert G\rvert}\sum_{g_{i}\in G}\chi_{j}(g_{i})\chi_{k}(g_{i})^{*}=\delta_{j,k},$ (86)
@@
As long as $G$ is a compact group, there will be unitary irreducible representations $U_{g_{i}}$ via the Peter-Weyl theorem DeitmarWeyl . Hence $U_{g_{i}}$ can be implemented in a quantum circuit for compact groups $G$. Using controlled gates $U_{g_{i}}$, in which the operation $U_{g_{i}}$ is applied only when the ancilla qubit is in the state $\lvert g_{i}\rangle$, then the operator $S_{\text{SELECT}}$ is successfully implemented.
@@ 4.2.1 Circuit Scaling
which is super-polynomial but sub-exponential Fulton1991RepresentationTA . This would be a significant improvement in scaling compared to $\mathcal{O}(n!)$. This allows a trade-off between total qubits and the ease of unitary preparation. Note that in this section, we introduced a very general framework, and improvements in circuit efficiency to implement $\hat{\chi}$ may well be possible for specific groups, representations, and data types.
4.2.2 Probabilistic Scaling
@@
4.3 Symmetry Invariant Encodings for Point Cloud Data
The previous subsection outlines how a quantum state can be projected to any irreducible representation of a finite group, or linear combination thereof. In this subsection, we highlight how certain irreducible representation subspaces may be useful for a specific type of data when encoded into a quantum state, namely point cloud data.
A point cloud is a collection of three-dimensional vectors (the points) that when viewed as a collective represent an image. Amongst other applications they are often associated with computer vision algorithms, as a primary method for performing three-dimensional imaging is the use of the Light Detection and Ranging (LiDAR) system, which produces point clouds as its data output raut2023endtoend . We can consider a point cloud $P$ as a set of three-dimensional vectors $P=\{\mathbf{p}_{1},\mathbf{p}_{2},...,\mathbf{p}_{n}\}$ that overall forms an image. We can consider a form of quantum encoding in which each individual point $\mathbf{p}_{i}$ is encoded into a quantum state $\lvert\mathbf{p}_{i}\rangle$ leading to a separable quantum state for the overall point cloud as
@@ 4.3 Symmetry Invariant Encodings for Point Cloud Data
4.3.1 Permutation Invariance
The ordering of the points within the set $P$ does not affect the overall image; however, in general a machine learning algorithm may output different results depending on the ordering of the points in the data input array. A solution to prevent this by utilising permutation symmetric encodings for point clouds was suggested in previous work heredge2023permutation where a quantum superposition of all possible permutations was used, leading to a permutation invariant quantum encoding such that
@@ 4.3.1 Permutation Invariance
providing the result proposed in Equation 102 and recreating the work of heredge2023permutation as a special case of a more general framework.
4.3.2 Rotational Invariance
@@
We consider irreducible representations of the group $S_{4}$, where the representations of group elements $\sigma\in S_{4}$ correspond to the SWAP gates $U_{\sigma}=\text{SWAP}_{\sigma}$ that permute the constituent points $\lvert\mathbf{p}_{i}\rangle$ accordingly. We are able to use this because Schur-Weyl duality (see Appendix B for a detailed discussion) provides a relation between irreducible finite-dimensional representations between the symmetric group and general linear group. This means that for the one-dimensional trivial representation of $SU(2)^{\otimes n}$ of multiplicity $m$, there will correspond an $m$-dimensional representation of $S_{n}$ covering the same subspace Kirby_2018 ; kirby_thesis ; ragone2023representation . The character table for $S_{4}$ is shown in Figure 2.
@@
If we utilise the irreducible representation projection procedure to project the state $\lvert\mathbf{p}_{1}\rangle\otimes\lvert\mathbf{p}_{2}\rangle\otimes\lvert\mathbf{p}_{3}\rangle\otimes\lvert\mathbf{p}_{4}\rangle$ to the subspace of irreducible representation $r=3$, with characters $\chi_{3}$, then this corresponds to projecting to the basis states Kirby_2018 ; kirby_thesis ; ragone2023representation
@@
When writing the action of $(U\otimes U\otimes U\otimes U)$ in the Schur basis, as shown in Figure 14, it can be seen that the above basis states $\lvert d_{1}\rangle,\lvert d_{2}\rangle$ are invariant under the action of $(U\otimes U\otimes U\otimes U)$ as they are in the subspace associated with the trivial representation of $SU(2)^{\otimes 4}$ kirby_thesis ; Kirby_2018 . We are able to reach this space through projecting in $S_{4}$ due to Schur-Weyl duality, since the irreducible subspaces of $S_{4}$ and $SU(2)^{\otimes 4}$ are simultaneously block diagonalised. This means that we have effectively projected the quantum encoded point cloud state to a subspace which is invariant with respect to three-dimensional rotations of the point cloud. This is a desirable symmetry for point clouds, as they are naturally invariant to three-dimensional rotations.
@@
Figure 10: The overlap between two identical Point Clouds that have been rotated about a random direction in three dimensions as the amount of rotation is increased, for the rotational invariant encoding and generic noninvariant encoding. To generate this plot, we randomly generated $n=4$ points on a sphere of radius $1$ to act as the point cloud. We then encoded the $(\theta,\phi)$ radial coordinates of each point into a qubit using the specification of Equation 104. This results in an initial state $\lvert P\rangle_{\text{init}}=\lvert\mathbf{p}_{1}\rangle\lvert\mathbf{p}_{2}\rangle\lvert\mathbf{p}_{3}\rangle\lvert\mathbf{p}_{4}\rangle$. We then repeat the procedure after the point cloud data is transformed by a three-dimensional rotation in random direction by some angle $\Theta$ to produce non-invariant state $\lvert P\rangle_{\Theta}$. To generate rotationally invariant states $\lvert P_{\text{rot}}\rangle$ we apply the result in Theorem 3 to project $\lvert P\rangle_{\Theta}$ states to the irreducible representation denoted by $r=3$ as specified in Table 2, which via Schur-Weyl duality lie in the trivial representation subspace of $SU(2)^{\otimes 4}$ and are hence rotationally invariant under our data encoding set-up. This means we now have rotationally invariant encoded states $\lvert P_{\text{rot}}\rangle_{\Theta}$ for all $\Theta$ values. The blue line plots the overlap $\langle{P}_{\text{init}}\rvert{P_{\text{rot}}}\rangle_{\Theta}$ and the red line plots the overlap $\langle{P}_{\text{init}}\rvert{P}\rangle_{\Theta}$ as the point cloud data rotation angle $\Theta$ is varied over $[0,\pi]$. As the overlap $\langle{P}_{\text{init}}\rvert{P_{\text{rot}}}\rangle_{\Theta}$ is constant we therefore numerically confirm that projection to the irreducible representation denoted by $r=3$ creates rotationally invariant states.
4.4 Symmetric Subspace Amplification
@@
The rotationally invariant encoding discussed previously, as well as previous work on permutation invariant encodings heredge2023permutation have been implemented in a binary fashion: the initial quantum state is either projected to a fully permutation invariant state or remains unchanged. Such an implementation significantly reduces the dimensionality of the encoding, yielding benefits in certain scenarios through improved generalisation of the model. This section explores the potential advantages of introducing a continuous spectrum of invariance with respect to some symmetry, rather than adhering to binary extremes, and investigates whether optimising this aspect as a hyperparameter could enhance classification performance in quantum machine learning (QML) models.
@@
Recent studies suggest that permutation equivariant variational circuits may be classically tractable under specific conditions anschuetz_efficient_2023 ; goh2023liealgebraic , indicating that a drastic reduction in dimensionality might increase the likelihood of classical simulatability. This raises concerns that fully projecting to a polynomially sized subspace could render algorithms susceptible to classical simulation. To address these issues, we propose a novel approach that involves partially amplifying the portion of the quantum state into a given invariant space. Specifically, we will consider amplifying the permutation invariant subspace, although any irreducible subspace could be chosen. The projection to this amplified space is governed by a hyperparameter $\alpha$ which directly influences the expressibility of the encoding. This method provides enhanced control over training performance and facilitates the tuning of the model to achieve optimal results. For data types with inherent permutation symmetry, such as point cloud data, the optimal global classification function must also exhibit permutation invariance. However, the specific outcome of a QML model depends on the circuit architecture, classification protocol, and inherent limitations of QML models, suggesting that a purely permutation invariant QML algorithm may not necessarily be the closest to the global solution. We hypothesise that by incrementally adjusting the degree of permutation invariance in the model, we can converge towards a value that is closer to the true global optimum performance than is possible with an non-invariant, or fully permutation invariant model. This concept is illustrated in Figure 11.
@@
Figure 11: Diagram providing a conceptual overview of the symmetric subspace amplification motivation. We start at $\alpha=0$ with a state encoding $\lvert P_{\text{start}}\rangle=\bigotimes_{i=1}^{n}\lvert\mathbf{p}_{i}\rangle$ corresponding to an initial state without symmetry amplification, which is input to some classification model $f(\cdot)$. As $\alpha$ increases, the state moves through the solution space until finally we have a fully symmetric state at $\alpha=1$ where the input state will be equal to $\lvert P_{\text{end}}\rangle=\frac{1}{\Omega^{\prime}}\sum_{\sigma\in S_{n}}\lvert\textbf{p}_{\sigma_{1}}\rangle\lvert\textbf{p}_{\sigma_{2}}\rangle...\lvert\textbf{p}_{\sigma_{n}}\rangle$ as initially proposed in heredge2023permutation . If we are dealing with data that is known to be permutation symmetric, then the global optimal model for the input data $\tilde{f}(\mathbf{x})_{\text{global}}$ (which could be classical or quantum) must be within the fully permutation symmetric subspace of models. However, the precise model that yields the global optimum remains unknown and may not correspond to the function $f(\cdot)$ used in this particular case, suggesting that an intermediate value of $\alpha$ might align more closely with the global optimum than the fully symmetric projection at $\alpha=1$. In the figure depicted, the red line illustrates the deviation between the global model solution and the nearest solution obtained at an intermediate value of $\alpha$ in which the input state is denoted $\lvert P_{\text{opt}}\rangle$
@@
In order to investigate the potential of symmetric subspace amplification, we implemented the model proposed here utilising the aforementioned technique alongside Qiskit statevector_simulator qiskit . This was used to encode point cloud data and perform classification on the sphere and torus point cloud dataset as specified in heredge2023permutation . More details can be found in Appendix C.
We record the average test accuracy over 10 experiments for a given $\alpha$ and then vary $\alpha$ to see which value is optimal for the data. The results presented in Figure 13 illustrate that increasing the degree of symmetry yields the highest test accuracy at an intermediate value, demonstrating the potential advantage of this technique. Moreover, it is observed that while a fully symmetric projection ($\alpha=1$) surpasses the performance of the non-symmetrised configuration ($\alpha=0$), there is a significant reduction in dimensionality as we approach a fully symmetric state. This reduction corresponds to projecting onto a polynomial size space for the case of symmetric subspaces heredge2023permutation . A key concern with fully symmetric states is their potential for efficient classical simulation due to this polynomially sized space. In contrast, the technique of partial symmetrisation not only achieves higher accuracy but also resides in a higher-dimensional space, possibly making it more challenging to simulate classically.
Figure 13: Test accuracy and effective dimension as the degree of permutation symmetry is varied. Result is the average over 10 repeated experiments, each with 100 data points each of the sphere and torus classification dataset using quantum kernel estimation followed by a Support Vector Machine classification havlicek_supervised_2019 ; heredge2023permutation ; Heredge21 . The effective dimension is found by explicitly calculating the higher dimensional encoding, performing principal component analysis, and calculating the amount of dimensions required to explain 95% of the variance of the dataset. We show that an intermediate value of permutation symmetry around $\alpha=0.7$ results in the best performance for this classification.
5 Conclusion

### v2

@@ IV Irreducible Subspace Projections
IV.1 Representation Theory Preliminaries
Following standard texts on group and representation theory Fulton and Harris (1991); Ragone et al. (2023) we provide a brief summary and introduction to the underlying mathematical framework utilised in this section.
@@ IV.1 Representation Theory Preliminaries
If a representation does contains a non-trivial subrepresentation, then it is called a reducible representation. Representations can often be decomposed into a direct sum of their irreducible representations, in which case they are called completely reducible. Maschke’s Theorem Maschke (1898) states that any finite-dimensional representation of a finite group $G$ over a field $F$ will be completely reducible (so long as the field characteristic does not divide the order of the group). It is also the case that every finite-dimensional representation of compact Lie groups is completely reducible Hall (2015). If a representation $(U,V)$ is completely reducible, then there exists a basis in which we can write $U_{g}$ as
@@
IV.2 Subspace Projection Circuit
We start by observing a result from the representation theory of finite groups which states Jean-Pierre
@@ IV.2 Subspace Projection Circuit
This is a known result in representation theory; for more details, see reference texts such as Fulton and Harris (1991); Jean-Pierre . In this section, we explore a method for implementing this projection on a quantum circuit using the Linear Combination of Unitaries (LCU) technique. Specifically, we demonstrate the practicality of realising the projection described in Theorem 3 on a quantum device. In the most comprehensive scenario, we establish that a quantum circuit can execute combinations of such $\hat{P}_{r}$ projections at the same time via the LCU method.
@@
Let $\lvert\psi\rangle\in V=(\mathbb{C}^{2})^{\otimes n}$ be a quantum state encoded on $n$ qubits. Let $U:G\rightarrow SU(2^{n})$ be a unitary representation of a finite group $G$, where the representatives for group elements $g_{i}\in G$ are denoted $U_{g_{i}}\in SU(2^{n})$ and are unitary operators that acts on $\lvert\psi\rangle$ and can be implemented in a quantum device. Utilising the Linear Combination of Unitaries framework on a quantum circuit, one can probabilistically apply a linear combination of irreducible representation projections
@@
where the $u_{i,j}$ will never affect the algorithm, but are required as the matrices involved must be of size $2^{k}$. The first $R$ columns correspond to the vectors that contain the characters for each element in a given representation. Considering the character orthogonality theorem Huppert (1998) which states that
$\frac{1}{\rvert G\rvert}\sum_{g_{i}\in G}\chi_{j}(g_{i})\chi_{k}(g_{i})^{*}=\delta_{jk},$ (91)
@@
As long as $G$ is a compact group there will be unitary irreducible representations $U_{g_{i}}$ via the Peter-Weyl theorem Deitmar (2002). Hence $U_{g_{i}}$ can be implemented in a quantum circuit for compact groups $G$. Using controlled gates $U_{g_{i}}$, in which the operation $U_{g_{i}}$ is applied only when the ancilla qubit is in the state $\lvert g_{i}\rangle$, then the operator $S_{\text{SELECT}}$ is successfully implemented.
@@ IV.2.1 Circuit Scaling
which is super-polynomial but sub-exponential Fulton and Harris (1991). This would be a significant improvement in scaling compared to $\mathcal{O}(n!)$. This allows a trade-off between total qubits and the ease of unitary preparation. Note that in this section, we introduced a very general framework, and improvements in circuit efficiency to implement $\hat{\chi}$ may well be possible for specific groups, representations, and data types.
IV.2.2 Probabilistic Scaling
@@
IV.3 Symmetry Invariant Encodings for Point Cloud Data
In this subsection we discuss projections to a single irreducible subspace in order to highlight how these subspaces can correspond to symmetries of the input data. However, it should be noted that these projections could take the quantum state to a polynomially sized space which may be classically simulatable, as has previously been shown to be the case with equivariant variational models Anschuetz et al. (2023); Goh et al. (2023). The following subsection will address this issue by instead considering all subspaces, and hence maintaining an exponentially large space overall, but allowing certain subspaces to be amplified relative to the others. Investigating single subspace projections remains useful in this context as it allows the identification of which irreducible representation subspace correspond to symmetries of the underlying data. We therefore proceed to highlight how certain irreducible representation subspaces correspond to symmetries of the underlying data when it has been encoded into a quantum state. In particular we focus on point cloud data due to its inherent permutation and rotation symmetry.
A point cloud is a collection of three-dimensional vectors (the points) that when viewed as a collective represent an image. Amongst other applications they are often associated with computer vision algorithms, as a primary method for performing three-dimensional imaging is the use of the Light Detection and Ranging (LiDAR) system, which produces point clouds as its data output Raut and Patole (2023). We can consider a point cloud $P$ as a set of three-dimensional vectors $P=\{\mathbf{p}_{1},\mathbf{p}_{2},...,\mathbf{p}_{n}\}$ that overall forms an image. We can consider a form of quantum encoding in which each individual point $\mathbf{p}_{i}$ is encoded into a quantum state $\lvert\mathbf{p}_{i}\rangle$ leading to a separable quantum state for the overall point cloud as
@@ IV.3 Symmetry Invariant Encodings for Point Cloud Data
IV.3.1 Permutation Invariance
The ordering of the points within the set $P$ does not affect the overall image; however, in general a machine learning algorithm may output different results depending on the ordering of the points in the data input array. A solution to prevent this by utilising permutation symmetric encodings for point clouds was suggested in previous work Heredge et al. (2024) where a quantum superposition of all possible permutations was used, leading to a permutation invariant quantum encoding such that
@@ IV.3.1 Permutation Invariance
providing the result proposed in Equation 107 and recreating the work of Heredge et al. (2024) as a special case of a more general framework.
IV.3.2 Rotational Invariance
@@
We consider irreducible representations of the group $S_{4}$, where the representations of group elements $\sigma\in S_{4}$ correspond to the SWAP gates $U_{\sigma}=\text{SWAP}_{\sigma}$ that permute the constituent points $\lvert\mathbf{p}_{i}\rangle$ accordingly. We are able to use this because Schur-Weyl duality (see Appendix B for a detailed discussion) provides a relation between irreducible finite-dimensional representations between the symmetric group and general linear group. This means that for the one-dimensional trivial representation of $SU(2)^{\otimes n}$ of multiplicity $m$, there will correspond an $m$-dimensional representation of $S_{n}$ covering the same subspace Kirby and Strauch (2018); Kirby ; Ragone et al. (2023). The character table for $S_{4}$ is shown in Figure 2.
@@
If we utilise the irreducible representation projection procedure to project the state $\lvert\mathbf{p}_{1}\rangle\otimes\lvert\mathbf{p}_{2}\rangle\otimes\lvert\mathbf{p}_{3}\rangle\otimes\lvert\mathbf{p}_{4}\rangle$ to the subspace of irreducible representation $r=3$, with characters $\chi_{3}$, then this corresponds to projecting to the basis states Kirby and Strauch (2018); Kirby ; Ragone et al. (2023)
@@
When writing the action of $(U\otimes U\otimes U\otimes U)$ in the Schur basis, as shown in Figure 21, it can be seen that the above basis states $\lvert d_{1}\rangle,\lvert d_{2}\rangle$ are invariant under the action of $(U\otimes U\otimes U\otimes U)$ as they are in the subspace associated with the trivial representation of $SU(2)^{\otimes 4}$ Kirby ; Kirby and Strauch (2018). We are able to reach this space through projecting in $S_{4}$ due to Schur-Weyl duality, since the irreducible subspaces of $S_{4}$ and $SU(2)^{\otimes 4}$ are simultaneously block diagonalised. This means that we have effectively projected the quantum encoded point cloud state to a subspace which is invariant with respect to three-dimensional rotations of the point cloud. This is a desirable symmetry for point clouds, as they are naturally invariant to three-dimensional rotations.
@@
Figure 17: The overlap between two identical Point Clouds that have been rotated about a random direction in three dimensions as the amount of rotation is increased, for the rotational invariant encoding and generic noninvariant encoding. To generate this plot, we randomly generated $n=4$ points on a sphere of radius $1$ to act as the point cloud. We then encoded the $(\theta,\phi)$ radial coordinates of each point into a qubit using the specification of Equation 109. This results in an initial state $\lvert P\rangle_{\text{init}}=\lvert\mathbf{p}_{1}\rangle\lvert\mathbf{p}_{2}\rangle\lvert\mathbf{p}_{3}\rangle\lvert\mathbf{p}_{4}\rangle$. We then repeat the procedure after the point cloud data is transformed by a three-dimensional rotation in random direction by some angle $\Theta$ to produce non-invariant state $\lvert P\rangle_{\Theta}$. To generate rotationally invariant states $\lvert P_{\text{rot}}\rangle$ we apply the result in Theorem 3 to project $\lvert P\rangle_{\Theta}$ states to the irreducible representation denoted by $r=3$ as specified in Table 2, which via Schur-Weyl duality lie in the trivial representation subspace of $SU(2)^{\otimes 4}$ and are hence rotationally invariant under our data encoding set-up. This means that we now have rotationally invariant encoded states $\lvert P_{\text{rot}}\rangle_{\Theta}$ for all $\Theta$ values. The blue line plots the overlap $\langle{P}_{\text{init}}\rvert{P_{\text{rot}}}\rangle_{\Theta}$ and the red line plots the overlap $\langle{P}_{\text{init}}\rvert{P}\rangle_{\Theta}$ as the point cloud data rotation angle $\Theta$ is varied over $[0,\pi]$. As the overlap $\langle{P}_{\text{init}}\rvert{P_{\text{rot}}}\rangle_{\Theta}$ is constant we therefore numerically confirm that projection to the irreducible representation denoted by $r=3$ creates rotationally invariant states.
IV.4 Symmetric Subspace Amplification
@@
The rotationally invariant encoding discussed previously, as well as previous work on permutation invariant encodings Heredge et al. (2024) have been implemented in a binary fashion: the initial quantum state is either projected to a fully permutation invariant state or remains unchanged. Such an implementation significantly reduces the dimensionality of the encoding, yielding benefits in certain scenarios through improved generalisation of the model. This section explores the potential advantages of introducing a continuous spectrum of invariance with respect to some symmetry, rather than adhering to binary extremes, and investigates whether optimising this aspect as a hyperparameter could enhance classification performance in quantum machine learning (QML) models.
@@
Recent studies suggest that permutation equivariant variational circuits may be classically tractable under specific conditions Anschuetz et al. (2023); Goh et al. (2023), indicating that a drastic reduction in dimensionality might increase the likelihood of classical simulatability. This raises concerns that fully projecting to a polynomially sized subspace could render algorithms susceptible to classical simulation. To address these issues, we propose a novel approach that involves partially amplifying the portion of the quantum state into a given invariant space. Specifically, we will consider amplifying the permutation invariant subspace, although any irreducible subspace could be chosen. Apart from edge cases, the quantum state will still in general be exponential in dimension but the permutation symmetric subspace will have a higher weighting compared to all other subspaces; the amount of permutation symmetry amplification is governed by a hyperparameter $\alpha$ which will therefore also affect the expressivity of the encoding. This method provides enhanced control over training performance and facilitates the tuning of the model to achieve optimal results. For data types with inherent permutation symmetry, such as point cloud data, the optimal global classification function must also exhibit permutation invariance. However, the specific outcome of a QML model depends on the circuit architecture, classification protocol, and inherent limitations of QML models, suggesting that a purely permutation invariant QML algorithm may not necessarily be the closest to the global solution. We hypothesise that by incrementally adjusting the degree of permutation invariance in the model, we can converge towards a value that is closer to the true global optimum performance than is possible with an non-invariant, or fully permutation invariant model. This concept is illustrated in Figure 18.
@@
Figure 18: Diagram providing a conceptual overview of the symmetric subspace amplification motivation. We start at $\alpha=0$ with a state encoding $\lvert P_{\text{start}}\rangle=\bigotimes_{i=1}^{n}\lvert\mathbf{p}_{i}\rangle$ corresponding to an initial state without symmetry amplification, which is input to some classification model $f(\cdot)$. As $\alpha$ increases, the state moves through the solution space until finally we have a fully symmetric state at $\alpha=1$ where the input state will be equal to $\lvert P_{\text{end}}\rangle=\frac{1}{\Omega^{\prime}}\sum_{\sigma\in S_{n}}\lvert\textbf{p}_{\sigma_{1}}\rangle\lvert\textbf{p}_{\sigma_{2}}\rangle...\lvert\textbf{p}_{\sigma_{n}}\rangle$ as initially proposed in Heredge et al. (2024). If we are dealing with data that is known to be permutation symmetric, then the global optimal model for the input data $\tilde{f}(\mathbf{x})_{\text{global}}$ (which could be classical or quantum) must be within the fully permutation symmetric subspace of models. However, the precise model that yields the global optimum remains unknown and may not correspond to the function $f(\cdot)$ used in this particular case, suggesting that an intermediate value of $\alpha$ might align more closely with the global optimum than the fully symmetric projection at $\alpha=1$. In the figure depicted, the red line illustrates the deviation between the global model solution and the nearest solution obtained at an intermediate value of $\alpha$ in which the input state is denoted $\lvert P_{\text{opt}}\rangle$
@@
In order to investigate the potential of symmetric subspace amplification, we implemented the model proposed here utilising the aforementioned technique alongside Qiskit statevector_simulator Abraham et al. (2019). This was used to encode point cloud data and perform classification on the sphere and torus point cloud dataset as specified in Heredge et al. (2024). More details can be found in Appendix C.
We record the average test accuracy over 10 experiments for a given $\alpha$ and then vary $\alpha$ to see which value is optimal for the data. The results presented in Figure 20 illustrate that increasing the degree of symmetry yields the highest test accuracy at an intermediate value, demonstrating the potential advantage of this technique. Moreover, it is observed that while a fully symmetric projection ($\alpha=1$) surpasses the performance of the non-symmetrised configuration ($\alpha=0$), there is a significant reduction in dimensionality as we approach a fully symmetric state. This reduction corresponds to projecting onto a polynomial size space for the case of symmetric subspaces Heredge et al. (2024). The $\alpha$ parameter is the ratio of the symmetric subspace to all other subspaces. In general one could select any relative weighting for all the irreducible subspaces, at the cost of having more parameters to optimise. We envisage that in practice a grid search would be used to find these hyperparameter values and that they would be optimised using a cross validation set, allowing final classification performance to be assessed on an unseen testing set.
A key concern with fully symmetric states is their potential for efficient classical simulation due to this polynomially sized space. In contrast, the technique of partial symmetrisation not only achieves higher accuracy but also resides in a higher-dimensional space, possibly making it more challenging to simulate classically. While computer vision data can consist of thousands of points, there are example datasets that may be more suitable in the near term, such as particle physics collision classification tasks in which there may be less than 10 particles (points) in a given event Heredge et al. (2021). This would likely be a nearer term goal for the application of point cloud specific algorithms, although the method described in this paper could be applied to many symmetries across various different data types.
Figure 20: Test accuracy and effective dimension as the degree of permutation symmetry is varied. Result is the average over 10 repeated experiments, each with 100 data points each of the sphere and torus classification dataset using quantum kernel estimation followed by a Support Vector Machine classification Havlíček et al. (2019); Heredge et al. (2024, 2021). The effective dimension is found by explicitly calculating the higher dimensional encoding, performing principal component analysis, and calculating the amount of dimensions required to explain 95% of the variance of the dataset. We show that an intermediate value of permutation symmetry around $\alpha=0.7$ results in the best performance for this classification.
V Conclusion